The area of triangular flag is 72 square centimeters. Its altitude is 2 centimeters longer than twice its base. Find the lengths of the altitude and the base.
Answer by nyc_function(2741)

We use A = 1/2 times base times height.

72 = 1/2 times x times (2x + 2)

72 = (x/2)(2x + 2)...This is the equation you need to find your answer. Can you finish?